# Accelerated Parallel Tempering via Neural Transports

### v3

@@ 6 Experiments
In this section, we evaluate our proposed accelerated parallel tempering and its three variants on a variety of targets. We defer all experimental details to Appendix˜D.
@@ 6.1 Comparison of Acceleration Methods
We evaluate the three variants of APT introduced in Section˜5: Flow-APT, CMCD-APT and Diff-APT, against the baseline of PT using the linear path with reference $\pi_{0}={\mathcal{N}}(0,\mathrm{I})$, on a 40-mode Gaussian mixture model in 10 dimensions (GMM-$10$) (Midgley et al., 2023). Table˜1 compares the performance of PT vs the APT variants for $N=6,10,30$ chains. We choose these numbers to represent: (1) the smallest $N$ where standard PT obtains round trips (2) the regime $N\approx\Lambda$ where PT is stable, and (3) the regime where $N\gg\Lambda$ and PT is close to optimal. All three APT variants substantially improve over classical PT, with increased round trips with $N$ and $K$. In the small $N$ regime (1) resulted in a 10x to 100x improvement and highlights the potential of APT for challenging problems where PT struggles and parallel chains are constrained. By contrast, when $N\gg\Lambda$ and the improvements over standard PT are less dramatic, they might not offset the cost of additional neural evaluations. Notably, CMCD with $K=5$ and $N=30$ matches the performance of the theoretical limit $T/(2+2\Lambda)\approx 5,349$ for standard PT when $N\to\infty$ as predicted by Theorem˜2.
[…]
@@
Method Neural Call $(\downarrow)$ $\hat{\Lambda}$ $(\downarrow)$ R $(\uparrow)$ CN-R $(\uparrow)$ R $(\uparrow)$ CN-R $(\uparrow)$ R $(\uparrow)$ CN-R $(\uparrow)$
@@
Figure 2: Round trip metrics for $K$-step Diff-APT ($K=1,2,5$) and Diff-PT using the true diffusion path, and Linear-PT targeting GMM-$d$ for $d=2,10,50,100$ when using 30 chains. (Left) Round trip rate against $d$. (Right) Compute-normalised round trip rate against $d$.
@@ 6.2 Scaling with Dimensions
To understand the theoretical performance of $K$-step APT for a fixed number of chains when we scale the dimension $d$, we use the fact that the path of distributions $(\pi^{\text{VP}}_{s})_{s\in[0,1]}$ induced by a VP diffusion process initialised at GMM-$d$ is analytically tractable, in order to run $K$-step Diff-APT with the true diffusion path. In Figure˜2, we compare the resulting round trip rate and compute-normalised round trip rate of the different algorithms when we fix the number of chains to 30. For all values of $d$, the round trip rate of Diff-APT monotonically increases over PT as we increase $K$, with greater gains for larger values of $d$, demonstrating that the accelerated swap improves the communication of states over the standard PT swap. We see a substantial improvement in round trip rate (right) with acceleration compared to PT ($K=0$) and a minor difference between algorithms as $K$ increases when normalised for compute (Right), suggesting the extra computation for APT is justified.
6.3 Free Energy Estimator
As we described in Section˜3.2, a by-product of Algorithm˜1 is the estimation of change in free energy $\Delta F$. We compare free energy estimates from CMCD-APT and Diff-APT against PT on two targets: DoubleWell(DW)-4, a particle system in Cartesian coordinates; and ManyWell(MW)-32, a highly multi-modal density introduced by Midgley et al. (2023). Figure˜3 presents box-plots of $30$ free energy estimates for DW-4 and MW-32 using 1,000 samples each.
Several key observations emerge: (1) Both CMCD-APT and Diff-APT exhibit markedly lower variance and bias than classical PT, across both targets. (2) For CMCD/Diff-APT, the variance and bias reduce steadily as $K$ increases. (3) For both PT and CMCD/Diff-APT, the variance and bias markedly decrease as the number of chains $N$ increases, thereby empirically validating Proposition˜1.
Figure 3: Estimates of $\Delta F$ for DW4 and ManyWell-32 by PT, CMCD-APT ($K=1,2,5$) and Diff-APT ($K=0,1,2,5$) using 1,000 samples. Each box consists of 30 estimates. The black dashed lines denotes the reference constant $\Delta F\approx 29.660$ estimated with PT using 60 chains and 100,000 samples and $\Delta F\approx 164.696$ from Midgley et al. (2023) for ManyWell-32.
@@ 6.4 Comparing APT with Neural Samplers
A significant advantage of APT is its asymptotic consistency: unlike most neural samplers, the Metropolis correction ensures APT does not incur a bias if the neural proposal is poorly trained. In Figure˜4, we demonstrate this property for both Diff-APT and CMCD-APT. Specifically, for CMCD-APT and Diff-APT, we take the trained model and map samples directly from the reference distribution to the target by concatenating the learned transports between each successive pair of intermediate distributions, labelled as CMCD and Diffusion respectively in Figure˜4.
As we can see, directly using the learned neural sampler dramatically drops performance, especially for small $K$. The diffusion model performs better than CMCD but still misallocates probability mass across two modes. On the contrary, all variants of CMCD-APT and Diff-APT recover the correct mode weights and align closely with the ground truth.
@@
Method $\hat{\Lambda}$ $(\downarrow)$ R $(\uparrow)$ CN-R $(\uparrow)$
@@ 6.5 Alanine Dipeptide
Finally, to provide a realistic target on which to evaluate APT, we take the Boltzmann distribution of Alanine Dipeptide in Cartesian coordinates at 300K, where we compare PT against CMCD-APT with $N=4$. This is a small molecule with 22 atoms (66 dimensions in total) which is highly challenging to sample from, due to many prohibitive energy barriers and high multimodality. Following the convention of molecular dynamics, we take the reference distribution to be a tempered version of the target, corresponding to 1200K. We run each algorithm for $T=50,000$ iterations and report the results in Table 2. We can see that CMCD-APT is able to provide significant acceleration in this realistic setting.
@@ 7 Conclusion
In this work, we formalise a framework integrating neural samplers to improve the sample efficiency of PT. However, several limitations, common to neural samplers, remain, including the additional burden of well optimising a neural network which has a large impact on performance - i.e. a poorly trained neural sampler may underperform PT. Further, while our approach accelerates PT by increasing the round trip rate, it incurs additional neural network evaluations. This can be computationally expensive for complex architectures. Future work should therefore develop principled criteria for deciding when to rely on PT or accelerated PT for robustness.
@@ Acknowledgements
Appendix A Theoretical Analysis of Accelerated PT
A.1 Further Details of Round Trips
Consider the APT Markov chain $\mathbf{X}_{t}=(X^{0}_{t},\ldots,X^{N}_{t})$ constructed by Algorithm˜1. Assume there are $m=0,\ldots,N$ parallel threads, each storing a component of $\mathbf{X}_{t}$ at time $t$. During the communication step, it is equivalent to swap indices rather than states. This has computational advantages in a distributed implementation of PT since swapping states across machines incurs a cost of $O(d)$ compared to the $O(1)$ cost of swapping indices. By tracking how the indices are shuffled, we can recover the PT state and also track the communication between reference and target through the dynamics of the permutated indices. See (Syed et al., 2022; Surjanovic et al., 2023, Algorithm 5) for details on implementing PT with a distributed implementation.
@@ A.1 Further Details of Round Trips
A.2 Proof of Theorem˜1
A.2.1 Proof of Ergodicity (Theorem˜1, Part 1)
@@
To arrive at Equation˜8 we need to show that
@@
Thus Equation˜9 is established.
Finally, the accelerated-PT Markov chain is aperiodic because the swaps can be rejected. The chain is clearly irreducible if each exploration kernel is irreducible. (In fact it can be proved, using more complicated arguments, that the chain is irreducible if the exploration kernel for the reference is irreducible and, for all $n$, the two distributions ${\mathbb{P}}_{K}^{n-1}$ and ${\mathbb{Q}}_{K}^{n}$ are mutually absolutely continuous.) By Roberts & Rosenthal (2004, Theorem 4, Fact 5), the Markov chain is ergodic and in particular the law of large numbers hold. ∎
A.2.2 Proof of Rejection Rate (Theorem˜1, Part 2)
@@
where $\mathbb{E}_{{\mathbb{P}}\otimes{\mathbb{Q}}}$ is a shorthand for the expectation under ${\mathbb{P}}^{n-1}_{K}\otimes{\mathbb{Q}}^{n-1}_{K}$ and the acceptance ratio $\alpha_{K}^{n}$ is given by
@@
From Equation˜10, Equation˜11, and Lemma˜1 below we have
@@
Let $\mu_{1}$ and $\mu_{2}$ be two mutually absolutely continuous measures. Then
@@
Combining Equation˜12 and Equation˜13 we get the desired result. ∎
A.2.3 Proof of Proposition˜1
@@
$\exp(-\Delta F_{n})\frac{\mathrm{d}{\mathbb{Q}}^{n}_{K}}{\mathrm{d}{\mathbb{P}}^{n-1}_{K}}=\exp(-W^{n}_{K}),\quad\exp(\Delta F_{n})\frac{\mathrm{d}{\mathbb{P}}^{n-1}_{K}}{\mathrm{d}{\mathbb{Q}}^{n}_{K}}=\exp(W^{n}_{K}).$
By taking expectations of the left and right expressions with respect to ${\mathbb{P}}^{n-1}_{K}$ and ${\mathbb{Q}}^{n}_{K}$ respectively, and taking the product over $n=1,\ldots,N$, we obtain the following expressions for $\exp(-\Delta F)$ and $\exp(\Delta F)$ respectively:
$\exp(-\Delta F)=\prod_{n=1}^{N}{\mathbb{P}}^{n-1}_{K}\left[\exp\left(-W^{n}_{K}\right)\right],\quad\exp(\Delta F)=\prod_{n=1}^{N}{\mathbb{Q}}^{n}_{K}\left[\exp\left(W^{n}_{K}\right)\right].$ (14)
Since $\textbf{X}_{t}$ is ergodic, $\exp(-\Delta\accentset{\rightharpoonup}{F}_{T})$ and $\exp(\Delta\accentset{\leftharpoonup}{F}_{T})$ are consistent estimators for $\exp(-\Delta F)$ and $\exp(\Delta F)$ respectively. Therefore $\Delta\hat{F}_{T}=\frac{1}{2}(\Delta\accentset{\rightharpoonup}{F}_{T}+\Delta\accentset{\leftharpoonup}{F}_{T})$ is consistent as well. Finally, note that if ${\mathbb{P}}^{n-1}_{K}={\mathbb{Q}}^{n}_{K}$, then $W^{n}_{K}(x_{0:K})\stackrel{{\scriptstyle a.s.}}{{=}}\Delta F_{n}$, and hence $\Delta\accentset{\rightharpoonup}{F}_{T}=\Delta\accentset{\leftharpoonup}{F}_{T}\stackrel{{\scriptstyle a.s.}}{{=}}\Delta F$. ∎
A.3 Accelerated PT as Vanilla PT on Extended Space
In particular we stress that the distributions $\pi_{\operatorname{ex}}^{n}$ are supported on $\mathcal{X}^{N}$ and not $\mathcal{X}$.
Proposition 4.
Proof.
Since the rejection rate only depends on the Randon-Nikodym derivative, it suffices to verify that
which is straightforward from Equation˜15. ∎
A.4 Parallelism versus Acceleration Time
The following proposition analyses these local rates for both methods. As in Section˜A.3, the main idea is to find a vanilla PT equivalent for both algorithms. We define
$\tau^{-1}_{\textrm{VA}}(\mu_{0},\ldots,\mu_{K}):=2+\sum_{k=1}^{K}\frac{r(\mu_{k-1},\mu_{k})}{1-r(\mu_{k-1},\mu_{k})}$
Proposition 5.
Define the sequence of distributions $({\mathbb{S}}_{k})_{k=0}^{K}$ as
Then the following equalities hold
$\tau_{\textrm{TA}}$ $=\tau_{\textrm{VA}}({\mathbb{S}}_{0},{\mathbb{S}}_{K})$ (16)
Proof.
This proposition shows that it is preferable to use the parallel-accelerated method, as the quantity in Equation˜17 is generally greater than that of Equation˜16 thanks to the effect of the bridge between ${\mathbb{S}}_{0}$ and ${\mathbb{S}}_{K}$. However, in practice the time-accelerated method consumes less memory and so might be more suitable in certain circumstances.
Appendix B Further Details on Analysis of PT
B.1 Proof of Proposition˜2

### v4

@@ 6 Experiments
In this section, we evaluate our proposed variants of APT on a variety of targets. We defer further experimental details to Appendix E.
@@ 6.1 Comparison of Acceleration Methods
Table 1: PT versus APT with different acceleration methods, targeting a 40-mode Gaussian Mixture model (GMM-10) target in 10 dimensions and standard Gaussian reference using $N=6,10,30$ parallel chains for $T=100,000$ iterations. For each method, we report the round trips (R), round trips per target evaluation, denoted as compute-normalised round trips (CN-R), the number of neural network evaluations per parallel chain every iteration (Neural Calls), and $\Lambda_{K}$ estimated using $N=30$ chains $(\hat{\Lambda}_{K})$ .
@@
Method Neural Calls $(\downarrow)$ $\hat{\Lambda}_{K}$ $(\downarrow)$ R $(\uparrow)$ CN-R $(\uparrow)$ R $(\uparrow)$ CN-R $(\uparrow)$ R $(\uparrow)$ CN-R $(\uparrow)$
@@
We evaluate the three variants of APT introduced in Section˜5: NF-APT, CMCD-APT and Diff-APT, against the baseline of classical PT using the geometric path with reference $\eta={\mathcal{N}}(0,\mathbf{I})$, on a 40-mode Gaussian mixture model in 10 dimensions (GMM-$10$) (Midgley et al., 2023). Table˜1 compares the performance of these methods in terms of round trips achieved and, in order to control for the additional computational cost required by APT, their compute-normalised round trips; for further details, see Appendix E.3. We choose to evaluate the different methods for $N=6,10,30$ chains to cover [^4] We refer to the discussion in Section 5 and 6.3 of Syed et al. (2024) for this classification of regimes.: (1) the smallest $N$ where PT is able to obtain round trips (2) the regime $N\approx\Lambda$ where PT is stable ($\Lambda$ is the global barrier for PT), and (3) the regime where $N\gg\Lambda$ and PT is close to optimal.
We see all three APT variants substantially improve over PT with increased round trips as $N$ and $K$ increase, and even when controlling for the additional computation required by APT. For instance, in the small $N$ regime, (1) resulted in a 10x to 100x improvement and highlights the potential of APT for challenging problems where PT struggles and parallel chains are constrained. By contrast, when $N\gg\Lambda$ and the improvements over standard PT are less dramatic, they might not offset the cost of additional neural evaluations. Notably, CMCD with $K=5$ and $N=30$ outperforms the theoretical limit $T/(2+2\Lambda)\approx 5,349$ of round trips for classical PT (when $N\to\infty$) predicted by Theorem˜2.
Figure 2: Round trip metrics for $K$-step Diff-APT ($K=1,2,5$) and Diff-PT using the true diffusion path, and PT targeting GMM-$d$ for $d=2,10,50,100$ when using 30 chains. (Left) Round trip rate against $d$. (Right) Compute-normalised round trip rate against $d$.
@@ 6.2 Scaling with Dimensions
To understand the theoretical performance of $K$-step APT for a fixed number of chains when we scale the dimension $d$, we use the fact that the path of distributions $(\pi^{\text{VP}}_{s})_{s\in[0,1]}$ induced by a VP diffusion process initialised at GMM-$d$ is analytically tractable, in order to run $K$-step Diff-APT with the true diffusion path. We compare this against the same PT setup as previously. In Figure˜2, we plot the resulting round trip rate and compute-normalised round trip rate of the different algorithms when we fix $N=30$; for further details, see Appendix E.4. For all values of $d$, the round trip rate of Diff-APT monotonically increases over PT as we increase $K$, with greater gains for larger values of $d$, demonstrating that the accelerated swap improves the communication of states over the standard PT swap. We see a substantial improvement in round trip rate (right) with acceleration compared to PT ($K=0$) and a minor difference between algorithms as $K$ increases when normalised for compute (Right), suggesting the extra computation for APT is justified.
6.3 Log-Normalising Constant (Free-Energy) Estimation
As we described in Section˜3.2, a by-product of Algorithm˜1 is the estimation of change in free energy $\Delta F=-\log Z$. We compare free energy estimates from CMCD-APT and Diff-APT against PT on two targets: DoubleWell-4 (DW-4), a particle system in Cartesian coordinates; and ManyWell-32 (MW-32), a highly multi-modal density introduced by Midgley et al. (2023). Figure˜3 presents box-plots of $30$ free energy estimates for DW-4 and MW-32 using 1,000 samples each; for further details, see Appendix E.5.
Figure 3: Estimates of $\Delta F$ for DW4 and ManyWell-32 by PT, CMCD-APT ($K=1,2,5$) and Diff-APT ($K=0,1,2,5$) using 1,000 samples. Each box consists of 30 estimates. The black dashed lines denote the reference constant $\Delta F\approx 29.660$ estimated with PT using 60 chains and 100,000 samples and $\Delta F\approx 164.696$ from Midgley et al. (2023) for ManyWell-32.
Several key observations emerge: (1) Both CMCD-APT and Diff-APT exhibit markedly lower variance and bias than PT across both targets. (2) For CMCD/Diff-APT, the variance and bias reduce steadily as $K$ increases. (3) For both PT and CMCD/Diff-APT, the variance and bias markedly decrease as the number of chains $N$ increases, thereby empirically validating Proposition˜1.
@@ 6.4 Comparing APT with Neural Samplers
A significant advantage of APT is its asymptotic consistency—unlike most neural samplers, the Metropolis correction ensures APT does not incur a bias if the neural transport is poorly trained. To demonstrate this advantage, for CMCD-APT and Diff-APT on DW-4, we take the trained forward transition kernels $P_{k}^{n-1}$ and map samples directly from the reference distribution to the target; we denote these respective new samplers by CMCD and Diffusion respectively and we compare the performance of these neural samplers to CMCD-APT and Diff-APT in Figure˜4; for further details, see Appendix E.6. To aid with visualisation, we plot the interatomic distances of samples in DW-4.
As we can see, directly using the learnt neural samplers dramatically drops performance, especially for small $K$. Out of the two, diffusion performs better than CMCD but still misallocates probability mass across two modes. On the contrary, all variants of CMCD-APT and Diff-APT recover the correct mode weights and align closely with the ground truth.
@@
Method $\hat{\Lambda}_{K}$ $(\downarrow)$ R $(\uparrow)$ CN-R $(\uparrow)$
@@ 6.5 Alanine Dipeptide
Finally, to provide a realistic target on which to evaluate APT, we take the Boltzmann distribution of Alanine Dipeptide in Cartesian coordinates at 300K, and we compare PT against our best performing APT variant: CMCD-APT with $N=4$. This is a small molecule with 22 atoms (66 dimensions in total) which is highly challenging to sample from due to many prohibitive energy barriers and high multimodality. Following the conventions of molecular dynamics, we take the reference distribution to be a tempered version of the target corresponding to 1200K. We run each algorithm for $T=50,000$ iterations and report the results in Table 2; for further details, see Appendix E.7. We can see that CMCD-APT is able to provide significant acceleration even in this difficult setting with increased round trips, as well as computed-normalised round trips, over PT.
@@ 7 Conclusion
In this work, we formalise the framework of Accelerated Parallel Tempering which integrates neural transports to improve the efficiency of PT. However, several limitations, common to neural samplers, remain, including the burden of well optimising a neural network (Grenioux & Noble, 2026) which has a large impact on performance—i.e. a poorly trained neural transport may result in underperformance. Further, while our approach accelerates PT by increasing the round trip rate, it incurs additional neural network evaluations. Future work should therefore develop principled criteria for deciding when to rely on PT or APT for robustness. On a different note, an interesting avenue for future work is developing APT for inference-time control of generative models (He et al., 2025c), following the similar application of SMC for these purposes.
@@ Acknowledgements
Appendix A Further Details on Round Trips
Consider the APT Markov chain $\mathbf{X}_{t}=(X^{0}_{t},\ldots,X^{N}_{t})$ constructed by Algorithm˜1. Assume there are $m=0,\ldots,N$ parallel threads, each storing a component of $\mathbf{X}_{t}$ at time $t$. During the communication step, it is equivalent to swap indices rather than states. This has computational advantages in a distributed implementation of PT since swapping states across machines incurs a cost of $O(d)$ compared to the $O(1)$ cost of swapping indices. By tracking how the indices are shuffled, we can recover the PT state and also track the communication between reference and target through the dynamics of the permuted indices. See (Syed et al., 2022; Surjanovic et al., 2023, Algorithm 5) for details on implementing PT with a distributed implementation.
@@ Appendix A Further Details on Round Trips
Appendix B Theoretical Analysis of Accelerated PT
B.1 Proof of Theorem˜1
B.1.1 Proof of Ergodicity (Theorem˜1, Part 1)
@@
To arrive at Equation˜6 we need to show that
@@
Thus Equation˜7 is established.
Finally, the accelerated-PT Markov chain is aperiodic because the swaps can be rejected. The chain is clearly irreducible if each exploration kernel is irreducible. (In fact it can be proved, using more complicated arguments, that the chain is irreducible if the exploration kernel for the reference is irreducible and, for all $n$, the two distributions ${\mathbb{P}}_{K}^{n-1}$ and ${\mathbb{Q}}_{K}^{n}$ are mutually absolutely continuous.) By Roberts & Rosenthal (2004, Theorem 4, Fact 5), the Markov chain is ergodic and in particular the law of large numbers holds. ∎
B.1.2 Proof of Rejection Rate (Theorem˜1, Part 2)
@@
where $\mathbb{E}_{{\mathbb{P}}\otimes{\mathbb{Q}}}$ is a shorthand for the expectation under ${\mathbb{P}}^{n-1}_{K}\otimes{\mathbb{Q}}^{n}_{K}$ and the acceptance ratio $\alpha_{K}^{n}$ is given by
@@
From Equation˜8, Equation˜9, and Lemma˜1 below we have
@@
Let $\mu_{1}$ and $\mu_{2}$ be two mutually absolutely continuous probability measures. Then
@@
Combining Equation˜10 and Equation˜11 we get the desired result. ∎
B.1.3 Proof of Proposition˜1
@@
$\frac{Z_{n}}{Z_{n-1}}\frac{\mathrm{d}{\mathbb{Q}}^{n}_{K}}{\mathrm{d}{\mathbb{P}}^{n-1}_{K}}=w^{n}_{K},\quad\frac{Z_{n-1}}{Z_{n}}\frac{\mathrm{d}{\mathbb{P}}^{n-1}_{K}}{\mathrm{d}{\mathbb{Q}}^{n}_{K}}=\frac{1}{w^{n}_{K}}.$
By taking expectations of the left and right expressions with respect to ${\mathbb{P}}^{n-1}_{K}$ and ${\mathbb{Q}}^{n}_{K}$ respectively, and taking the product over $n=1,\ldots,N$, we obtain the following expressions for $Z$ and $Z^{-1}$ respectively:
$Z=\prod_{n=1}^{N}{\mathbb{P}}^{n-1}_{K}\left[w^{n}_{K}\right],\quad Z^{-1}=\prod_{n=1}^{N}{\mathbb{Q}}^{n}_{K}\left[(w^{n}_{K})^{-1}\right].$ (12)
Since $\textbf{X}_{t}$ is ergodic, $\accentset{\rightharpoonup}{Z}_{T}$ and $\accentset{\leftharpoonup}{Z}_{T}$ are consistent estimators for $Z$ respectively. Therefore $\hat{Z}_{T}=(\accentset{\rightharpoonup}{Z}_{T}\accentset{\leftharpoonup}{Z}_{T})^{1/2}$ is consistent as well. Finally, note that if ${\mathbb{P}}^{n-1}_{K}={\mathbb{Q}}^{n}_{K}$, then $w^{n}_{K}(x_{0:K})\stackrel{{\scriptstyle a.s.}}{{=}}Z_{n}/Z_{n-1}$, and hence $\accentset{\rightharpoonup}{Z}_{T}=\accentset{\leftharpoonup}{Z}_{T}\stackrel{{\scriptstyle a.s.}}{{=}}Z$. ∎
B.2 Proof of Proposition˜2